### v1

@@ 8 Summary
  • The clusters showed tidal tails extending beyond their Jacobi radius, ranging from one to four times the Jacobi radius.
@@
  • The M67 cluster and its tidal tails displayed a sign of rotation in the sky plane, aligning with predictions from $N$-body simulations and confirmed through rotational analysis, indicating that the extended structure is indeed a tidal tail. While NGC 2281 also revealed rotational features after the rotational analysis, these were confined primarily to their central cluster regions. For the remaining clusters, the available radial velocity data were insufficient to confidently assess the presence of rotation.
  • In this study, the binary fraction and the luminosity function of the tails and the main cluster are compared. In general, the binary fraction in the tidal tails is greater than in the cluster. Concerning the mass function, we observed that the overall slope remained consistent between clusters and tidal tails across all five clusters, but all tails showed a deficiency of high-mass stars.
Our results showed that carefully optimized clustering methods can identify extended structures around star clusters up to a kiloparsec. This broadens the range of clusters that can be studied compared to methods relying solely on projected on-sky parameters (e.g., Bhattacharya et al. 2022; Tarricq et al. 2022), probabilistic approaches (e.g., Kos 2024), or techniques constrained by astrometric precision, such as the convergent point method, which is typically limited to within $\approx$400 pc (Risbud et al. 2025). Detecting rotational patterns in outer regions also serves as an important tool for validating the presence of tidal tails. Extending this analysis to a larger sample of clusters could reveal hidden tidal structures in the Galactic field and provide important observational constraints for theoretical analysis.
@@
The catalogue of cluster and tidal tail members is available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/.
@@ Acknowledgements.
This work acknowledges the valuable guidance of Prof. Kulinder Pal Singh, who provided critical support during the initial stages of this research. VJ thanks the Alexander von Humboldt Foundation for their support. This study has used data from the European Space Agency (ESA) Gaia mission (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement.
@@ A.2 Appendix Figures of individual clusters and their description
The caption to the Appendix Figs. below 7 to 11 is as follows: (a) Spatial distribution in X-Y coordinates (b) Spatial distribution in Y-Z coordinates. (c) Spatial distribution in X-Z coordinates. (d) Spatial distribution in equatorial coordinates. (e) Apparent CMD. (f) Absolute CMD. (g) Vector point diagram. (h) LF (see Subsection 3.6) for the cluster members (blue dots) and tails (coral dots).
Figure 7: Results from our analysis for BH 164: The image description can be found in the Appendix A.2.
Figure 8: Same as the Figure 7 for Alessi 2.
Figure 9: Same as the Figure 7 for NGC 2281.
Figure 10: Same as the Figure 7 for NGC 2354.
Figure 11: Same as the Figure 7 for M67.
Figure 12: NGC 2281 Rotation detailed in Subsection 6.4. The red dotted line is the best-fitting Plummer model profile, obtained using a core radius of 1.5 pc, fitted to the observed RMS rotational velocities.

### v3

@@ 8 Summary
  • The clusters showed tidal tails extending beyond their Jacobi limit, ranging from one to four times their $r_{J}$.
@@
  • The M67 cluster and its tidal tails displayed a sign of rotation in the sky plane, aligning with predictions from N-body simulations and confirmed through rotational analysis, indicating that the extended structure is indeed a tidal tail. While NGC 2281 also revealed rotational features after the rotational analysis, these were confined primarily to its central cluster regions. For the remaining clusters, the available RV data were insufficient to confidently assess the presence of rotation.
  • In this study, the binary fraction and the LF of the tails and the main cluster are compared. In general, the binary fraction in the tidal tails is greater than in the cluster. Concerning the mass function, we observed that the overall slope remained consistent between clusters and tidal tails across all five clusters, but all tails showed a deficiency of high-mass stars.
Our results showed that carefully optimized clustering methods can identify extended structures around star clusters up to a kiloparsec. This broadens the range of clusters that can be studied compared to methods relying solely on projected on-sky parameters (e.g., Bhattacharya et al. 2022; Tarricq et al. 2022), probabilistic approaches (Kos 2024), or techniques constrained by astrometric precision, such as the convergent point method, which is typically limited to within $\approx$400 pc (Risbud et al. 2025). Detecting rotational patterns in outer regions also serves as an important tool for validating the presence of tidal tails. Extending this analysis to a larger sample of clusters could reveal hidden tidal structures in the Galactic field and provide important observational constraints for theoretical analysis.
@@
Table members.dat is only available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/.
@@ Acknowledgements.
We thank the anonymous referee for their constructive comments, which helped improve the manuscript. IS thanks Prof. Kulinder Pal Singh, who provided support during the initial stages of this research. VJ thanks the Alexander von Humboldt Foundation for their support. This study has used data from the European Space Agency (ESA) Gaia mission (https://www.cosmos.esa.int/Gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/Gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement.
@@ A.2 Appendix figures of individual clusters and their description
The caption to the Appendix Figs. below 7 to 11 is as follows: (a) Spatial distribution in X-Y coordinates (b) Spatial distribution in Y-Z coordinates. (c) Spatial distribution in X-Z coordinates. (d) Spatial distribution in equatorial coordinates. (e) Apparent CMD. (f) Absolute CMD. (g) Vector point diagram. (h) LF (see Subsect. 4.3) for the cluster members (blue dots) and tails (coral dots).
Figure 7: Results from our analysis for BH 164: Image description can be found in the Appendix A.2.
Figure 8: Same as the Fig. 7 for Alessi 2.
Figure 9: Same as the Fig. 7 for NGC 2281.
Figure 10: Same as the Fig. 7 for NGC 2354.
Figure 11: Same as the Fig. 7 for M67.
Figure 12: NGC 2281 rotation detailed in Subsect. 6.4. Red dotted line is the best-fitting Plummer model profile, obtained using a core radius of 1.5 pc, fitted to the observed RMS rotational velocities.